OpenStudy (anonymous):

help please ! In the xy. planem the graph of y=x^2 and the circle with center (0,1) and radius 3 hasve now how many points of intersection

OpenStudy (anonymous):

x^2+(y-1)^2 = 9 y+y^2-2y-8 = 0 y^2-y-8 = 0 Using the quadratic formula y = 3.373 (must be positive) x = 1.837 and -1.837 Intersection at 2 points (1.837, 3.373) and (-1.837, 3.373)

OpenStudy (anonymous):

How I use the quadratic formula in this?